Misty wishes to obtain 85 oz of a 40% solution by combining a 72% acid solution with a 25% acid solution. How much of each solution should Misty use.
I have gotten the equation of ".72x+.25(85-x)=34" but am not sure if that is correct. I also need to illustrate it so if you have any ideas, i would appreciate the help.

Answer by jorel1380(3719)

85x.4=34
m+n=85
.72m+.25n=35
72m+25n=3400
25m+25n=2125
47m=1275
m=1275/47=27.13
n=57.87
Misty needs 27.13 oz. of 72% acid and 57.87 oz. of 25% solution..